OpenStudy (anonymous):

Can someone please help me on this question, Let f(x) = 3x^2 - 6x, Evaluate and simplify of the expression f(t+h)

OpenStudy (saifoo.khan):

3x(x - 2)

OpenStudy (anonymous):

OH my bad haha

OpenStudy (saifoo.khan):

lol

OpenStudy (anonymous):

actually the answer is 3t^2 +6ht + 3h^2-6t-6h

OpenStudy (anonymous):

\[3(x+h)^2-6(x+h)=3(x^2+2hx+h^2)-6x-6h\]\[=3x^2+6xh+3h^2-6x-6h\]

OpenStudy (anonymous):

Oh so thats how you do it.. thanks a lot I appreciate it.